OpenStudy (anonymous):

Which conic section does the equation below describe (y-1)^2=8(x+3) circle, ellipse, parabola, hyperbola

OpenStudy (danjs):

work it all out and put it in a standard form, of x^2 +- y^2 = 1

OpenStudy (anonymous):

would it be y^2+1=8x+24?

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

hi

OpenStudy (anonymous):

is this right y^2+1=8x+24

OpenStudy (danjs):

Parabola with a horizontal axis. Vertex at (h,k) Focus is (h + p, k). Directrix is the line x = h – p. Axis is the line y = k. (y – k)^2 = 4p(x – h) and \[p \neq 0\]

OpenStudy (danjs):

reload page

OpenStudy (anonymous):

how did you know that it was a parabola? I have 4 more questions like that and i would like to be able to know how to do them

OpenStudy (danjs):

a basic parabola takes the form y = x^2

OpenStudy (danjs):

or x = y^2

OpenStudy (anonymous):

for (x-1)^2/20-(y+2)^2/16=1 would it be a ellipse

OpenStudy (danjs):

here is a good link that will help you identify conic sections. Put the equation they give you into a standard form. http://hotmath.com/hotmath_help/topics/conic-sections-and-standard-forms-of-equations.html

OpenStudy (danjs):

no, an ellipse is x^2/a + y^2/b = 1 a Hyperbola has a minus sign instead of a plus

OpenStudy (danjs):

remember a circle is x^2 + y^2 = 1 an ellipse is just another less perfect kind of circle, that is what the 'a' and 'b' come in.

OpenStudy (anonymous):

so the answer to that problem would be a hyperbola?

OpenStudy (danjs):

yes, if it was a plus then it would be a ellipse

OpenStudy (anonymous):

(x+2)^2/16+(y-9)^2/36 is an ellipse, right?

OpenStudy (danjs):

what do you think?

OpenStudy (anonymous):

yes because it has a +

OpenStudy (danjs):

yep, look at the third entry on the table in that link

OpenStudy (danjs):

(h,k) =( -2,9)

OpenStudy (anonymous):

(x+2)^2=4(y-3) is a circle and 2x^2+2y^2-6x+4y+1=0 is a parabola

OpenStudy (danjs):

Whatever is in the parenthesis like (x-4)^2 is a translation of the center of the ellipse, and the divisor is what stretches the thing out

OpenStudy (danjs):

ok the first one is not a circle

OpenStudy (danjs):

a circle would contain, squares on both the x and the y terms

OpenStudy (danjs):

circle: x^2 + y^2 = 1 center at the origin and radius 1

OpenStudy (anonymous):

ugh. this is so confusing... the first one would be a parabola

OpenStudy (danjs):

yes, it is just like the first question you asked with different numbers.

OpenStudy (anonymous):

but the first question you said it was a circle

OpenStudy (danjs):

no, the original question you asked is a parbola.

OpenStudy (danjs):

If you only have to choose between parabola, hyperbola, and ellipse/circle just remember, a parabola only has one of the terms squared an ellipse/circle has both positive x^2 and y^2 a hyperbola has one subtracted, a minus x^2 or y^2 term

OpenStudy (danjs):

For that one, Have you done "completing the square" in class? 2x^2+2y^2-6x+4y+1=0

OpenStudy (anonymous):

1) (y-1)^2=8(x+3) parabola 2) (x-1)^2/20 - (y+2)^2/16=1 hyperbola 3) (x+2)^2/16 + (y-9)^2/36=1 ellipse 4) (x+2)^2=4(y-3) parabola 5) 2x^2+2y^2-6x+4y+1=0 circle are these answers correct?

OpenStudy (danjs):

yes

OpenStudy (danjs):

to put the last one in a more familiar standard form for a circle, you need to "complete the square"

OpenStudy (danjs):

(x-h)^2 + (y-k)^2 = r^2 ---circle centered at (h,k) -- radius = r

OpenStudy (anonymous):

sorry my internet went out. this is my first time here on this site.. but I can see that I should like give a medal but i don't know how to do that for you or can I since im new

OpenStudy (danjs):

ha, i dont really know, just been here a couple days myself, maybe just hit best response button

OpenStudy (anonymous):

alright well thank you for you help!